OpenStudy (anonymous):

What is the factored form of the following expressions? x^2 – 10xy + 24y^2

OpenStudy (hoblos):

is it 24y^2 or 25y^2 ?? check it please

OpenStudy (anonymous):

yes i think so

OpenStudy (anonymous):

(x + 6y)(x + 4y) (x – 2y)(x + 12y) (x + 2y)(x – 12y) (x – 6y)(x – 4y)

OpenStudy (hoblos):

(x – 6y)(x – 4y)

OpenStudy (hoblos):

d

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so D?

OpenStudy (hoblos):

yes it is D

OpenStudy (anonymous):

okay, thanks!

OpenStudy (hoblos):

you are welcome :)

OpenStudy (anonymous):

\[(a+b)(a+c)=a ^{2}+(b+c)a+bc\]

OpenStudy (anonymous):

you can write this in any form (a+b)(a+c)=a^2 +ab+ac+bc